# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v2

@@ 1 Introduction
In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2020; Cohen et al., 2021, 2022). Precisely, if the step size is $\eta>0$, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations. Cohen et al. (2021) named this regime Edge of Stability (EoS). Surprisingly, unlike for quadratics, where this would imply divergence, this instability does not damage convergence when training neural networks.
@@
In the case of mini-batch algorithms, (i) Jastrzębski et al. (2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when. Moreover, Cohen et al. (2021) noticed that SGD somehow acclimates to the hyperparameters. However, they did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 1, often without a proper progressive sharpening phase. This scenario leaves the most basic questions open: In what way the location of convergence of SGD acclimates to the choice of hyperparameters? What are the key quantities involved? To be more specific, can we characterize the training phenomena in (i), (ii), (iii) above? What determines them? Does SGD train in an unstable regime?
@@ 2.1 Progressive Sharpening and Edge of Stability
Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018; Fort & Ganguli, 2019). Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training (Keskar et al., 2016; Sagun et al., 2016; Fort & Scherlis, 2019; Jastrzębski et al., 2019). Later, Jastrzębski et al. (2020) and Cohen et al. (2021) managed to make this precise, noticing that, along the trajectories of SGD and GD, $\lambda_{\max}$ increases steadily, often after a small number of steps of decrease. This phenomenon was called progressive sharpening by Cohen et al. (2021).
@@
Early studies (Goodfellow et al., 2016; Li et al., 2019; Jiang et al., 2019; Lewkowycz et al., 2020) revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction. Jastrzębski et al. (2020) attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening. This transition slows the convergence by confining the dynamics to a region of lower sharpness. Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape. Jastrzębski et al. (2020); Cohen et al. (2021, 2022) indeed showed that the phase transition comes at different points for different algorithms on the same landscapes. Jastrzębski et al. (2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it. Cohen et al. (2021, 2022) later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms.
@@
After the phase transition above, GD and full-batch train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value. The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics. Observations from Cohen et al. (2021, 2022) indicate that most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution. Lee & Jang (2023) explained why $\lambda_{\max}$ in this regime is generally above $2/\eta$, rarely aligning exactly. The reasons include: (i) nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and (ii) the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
@@ 2.2 SGD, Hessian, and Generalization
This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1994, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016). Analogously, recent work strongly aligns with our result: Jastrzębski et al. (2021) demonstrates that large learning rates in SGD act similarly to penalizing the trace of the Fisher matrix–which approximates the Hessian during training and shares its top eigenspaces (Jastrzębski et al., 2018; Martens, 2020; Thomas et al., 2020)–in image classification tasks.
@@
SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance (LeCun et al., 2012; Keskar et al., 2016; Goyal et al., 2017; Jastrzębski et al., 2018; Masters & Luschi, 2018; Smith et al., 2021). This advantage has been widely attributed to the flatness of the minima (Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020). Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks (Izmailov et al., 2019; Foret et al., 2021). Moreover, Jastrzębski et al. (2021) shows that, in practice, penalizing the trace of the Fisher information matrix empirically improves generalization and reduces memorization—especially when training with noisy labels. We believe that our quantification and the introduction of a theoretical framework for these improvements in sharpness may be a stepping stone towards understanding these generalization benefits theoretically.
@@ 3 SGD Typically Occurs at the EoSS
Figure 3: Comparing different sharpness measures. Red: observed Batch Sharpness on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$. Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
@@ 3. Catapults:
Unlike EoS, EoSS stabilizes the expectation of a quantity which the algorithm sees one observation at time. Occasionally, a series of batches sampled have outlying sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 4). This is typically followed by renewed progressive sharpening, returning to the EoSS regime. This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., (Lewkowycz et al., 2020; Zhu et al., 2024).
@@ 5.2 Size of $\lambda_{\max}$
Unfortunately, we could not find a way to relate the size of $\lambda_{\max}$ to learning rate and batch size–see Figure 9–if not through the gap it has with the average $\lambda_{\max}$ of the mini-batch Hessian $\lambda_{\max}^{b}$. Notably, the critical batch size we mention is the point beyond which training enters the full-batch regime. Works as (Zhang et al., 2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation". We conjecture there may be a relation between the two quantities and we leave it to future work.
@@ A.3 Formulating Stability for Mini-Batch SGD
Unlike in the deterministic case, we have to be a little more careful with the formulation here. Let $(\lambda_{i},v_{i})$ be the ordered eigenvalue-eigenvector pairs of $\mathbf{A}$. With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see (Cohen et al., 2021). The idea is that if a component $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ exists, after $k$ steps it gets multiplied by $(1-\eta\alpha_{i})^{k}$. This implies that the system will diverge if and only if $|1-\eta\alpha_{i}|>1$. When our step is a noise-injected GD similar stability phenomena happens, $\mathbf{A}$ is not impacted by the noise, see Appendix D.
@@
Thus the stability of (Cohen et al., 2021), from this perspective, is equivalent to enforcing for all $t$ the inequality Equation (9), or analogously enforcing it in expectation.
@@ Appendix B Stability for Mini-Batch SGD
Appendix C Comparison with (Lee & Jang, 2023)
After we wrote the first version of this article, we realized that a (Lee & Jang, 2023) already dealt with this problem. We draw here a comparison and show that they come to a different conclusion because they deal with a different notion of stability. While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of (Cohen et al., 2021).
@@ Appendix C Comparison with (Lee & Jang, 2023)
While the inequalities seem analogous and both equally motivated at first glance, we claim they carry an important difference. Enforcing Equation (12) does imply that in 1 step the loss does not go up. However, it does not imply that the loss will not go up eventually. The instability threshold discussed in (Cohen et al., 2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.
@@
However, after $T$ steps we can lower bound the loss with an increasing sequence which eventually diverges. Moreover, the loss asymptotically will approach the sequence with a linear convergence rate. Precisely, define the noise $\zeta_{B}$ such that $\nabla f_{B}=\nabla f+\zeta_{B}$. Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of (Lee & Jang, 2023) we may have decrease in loss in one step, indeed
@@ Appendix D Stability for Noise-Injected Gradients Descents
This imply that the Hessian does not change from step to step. The expected stability formulation is actually the one of (Lee & Jang, 2023) in the sense of (Cohen et al., 2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
@@ Appendix E Sampling the Mini Batches
In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of (Beneventano, 2023) as the tools previously developed by (Mishchenko et al., 2020; Smith et al., 2021) only work for infinitesimal step size and the tools in Appendix E of (Beneventano, 2023) are the only ones which allow us to consider more products and big learning rate. Let $k=n/b$ be the number of batches in an epoch and call $B_{j}$ the batch of $j-th$ step and $\mathbf{a}_{B_{j}},\mathbf{b}_{B_{j}}$ the averages of $\mathbf{a}$ and $\mathbf{b}$ over the batch $B_{j}$.
@@
Here if $\eta=O\left(\frac{b^{2}}{n}[\mathrm{Var}(\mathbf{a}_{1})]^{-1}\right)$ which is generally very small, the second them may already be of size $O(1)$. We already see that in this case $\eta$ of the instability threshold has to be set also through the variance of the $\mathbf{a}_{i}$, the size of the batch or the number of batches. More generally, by applying Proposition ? of (Beneventano, 2023) we obtain that

### v3

@@ 1 Introduction
In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2019, 2020; Cohen et al., 2021, 2022). Precisely, if the step size is $\eta>0$ for full-batch GD, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations. Cohen et al. (2021) named this regime Edge of Stability (EoS). Surprisingly, unlike for quadratics, where this would imply divergence, this instability does not damage convergence when training neural networks.
@@
In the case of mini-batch algorithms, (i) Jastrzębski et al. (2019, 2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when. Moreover, Cohen et al. (2021) noticed that SGD somehow acclimates to the hyperparameters. However, they did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 1, often without a proper progressive sharpening phase. This scenario leaves the most basic questions open: In what way the location of convergence of SGD acclimates to the choice of hyperparameters? What are the key quantities involved? To be more specific, can we characterize the training phenomena in (i), (ii), (iii) above? What determines them? Does SGD train in an unstable regime?
@@ 2.1 Progressive Sharpening and Edge of Stability
Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018, 2019; Fort & Ganguli, 2019). Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training (Keskar et al., 2016; Sagun et al., 2016; Fort & Scherlis, 2019; Jastrzębski et al., 2019). Later, Jastrzębski et al. (2019, 2020) and Cohen et al. (2021) managed to make this precise, noticing that, along the trajectories of SGD and GD, $\lambda_{\max}$ increases steadily, often after a small number of steps of decrease. This phenomenon was called progressive sharpening by Cohen et al. (2021).
@@
Early studies (Goodfellow et al., 2016; Li et al., 2019; Jiang et al., 2019; Lewkowycz et al., 2020) revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction. Jastrzębski et al. (2020) attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening. This transition slows the convergence by confining the dynamics to a region of lower sharpness. Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape. Jastrzębski et al. (2019, 2020); Cohen et al. (2021, 2022) indeed showed that the phase transition comes at different points for different algorithms on the same landscapes. Jastrzębski et al. (2019, 2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it. Cohen et al. (2021, 2022) later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms.
@@
After the phase transition above, GD and full-batch Adam train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value. The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics. Observations from Cohen et al. (2021, 2022) indicate that most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution. Lee & Jang (2023) explained why $\lambda_{\max}$ in this regime is generally above $2/\eta$, rarely aligning exactly. The reasons include: (i) nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and (ii) the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
@@ 2.2 SGD, Hessian, and Generalization
This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016). Analogously, recent work strongly aligns with our result: Jastrzębski et al. (2021) demonstrates that large learning rates in SGD act similarly to penalizing the trace of the Fisher matrix–which approximates the Hessian during training and shares its top eigenspaces (Jastrzębski et al., 2018; Martens, 2020; Thomas et al., 2020)–in image classification tasks.
@@
SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance (LeCun et al., 2012; Keskar et al., 2016; Goyal et al., 2017; Jastrzębski et al., 2018; Masters & Luschi, 2018; Smith et al., 2021). This advantage has been widely attributed to the flatness of the minima (Murata et al., 1994; Hochreiter & Schmidhuber, 1994; Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020). Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks (Izmailov et al., 2019; Foret et al., 2021). Moreover, Jastrzębski et al. (2021) shows that, in practice, penalizing the trace of the Fisher information matrix empirically improves generalization and reduces memorization—especially when training with noisy labels. We believe that our quantification and the introduction of a theoretical framework for these improvements in sharpness may be a stepping stone towards understanding these generalization benefits theoretically.
@@ 3 SGD Typically Occurs at the EoSS
Figure 3: Comparing different sharpness measures. Red: observed sharpness on the step’s mini batch on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$. Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
@@ 3. Catapults:
Unlike EoS, EoSS stabilizes the expectation of a quantity which the algorithm sees one observation at time. Occasionally, a series of batches sampled have outlying sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 4). This is typically followed by renewed progressive sharpening, returning to the EoSS regime. This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., Lewkowycz et al. (2020); Zhu et al. (2024).
@@ 5.2 Size of $\lambda_{\max}$
Unfortunately, we could not find a way to relate the size of $\lambda_{\max}$ to learning rate and batch size–see Figure 9–if not through the gap it has with the average $\lambda_{\max}$ of the mini-batch Hessian $\lambda_{\max}^{b}$. Notably, the critical batch size we mention is the point beyond which training enters the full-batch regime. Works as Zhang et al. (2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation". We conjecture there may be a relation between the two quantities and we leave it to future work.
@@ A.3 Formulating Stability for Mini-Batch SGD
Unlike in the deterministic case, we have to be a little more careful with the formulation here. Let $(\lambda_{i},v_{i})$ be the ordered eigenvalue-eigenvector pairs of $\mathbf{A}$. With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see Cohen et al. (2021). The idea is that if a component $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ exists, after $k$ steps it gets multiplied by $(1-\eta\alpha_{i})^{k}$. This implies that the system will diverge if and only if $|1-\eta\alpha_{i}|>1$. When our step is a noise-injected GD similar stability phenomena happens, $\mathbf{A}$ is not impacted by the noise, see Appendix D.
@@
Thus the stability of Cohen et al. (2021), from this perspective, is equivalent to enforcing for all $t$ the inequality Equation (9), or analogously enforcing it in expectation.
@@ Appendix B Stability for Mini-Batch SGD
Appendix C Comparison with Lee & Jang (2023)
After we wrote the first version of this article, we realized that Lee & Jang (2023) already dealt with this problem. We draw here a comparison and show that they come to a different conclusion because they deal with a different notion of stability. While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of Cohen et al. (2021).
@@ Appendix C Comparison with Lee & Jang (2023)
While the inequalities seem analogous and both equally motivated at first glance, we claim they carry an important difference. Enforcing Equation (12) does imply that in 1 step the loss does not go up. However, it does not imply that the loss will not go up eventually. The instability threshold discussed in Cohen et al. (2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.
@@
However, after $T$ steps we can lower bound the loss with an increasing sequence which eventually diverges. Moreover, the loss asymptotically will approach the sequence with a linear convergence rate. Precisely, define the noise $\zeta_{B}$ such that $\nabla f_{B}=\nabla f+\zeta_{B}$. Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of Lee & Jang (2023) we may have decrease in loss in one step, indeed
@@ Appendix D Stability for Noise-Injected Gradients Descents
This imply that the Hessian does not change from step to step. The expected stability formulation is actually the one of Lee & Jang (2023) in the sense of Cohen et al. (2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
@@ Appendix E Sampling the Mini Batches
In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of Beneventano (2023) as the tools previously developed by Mishchenko et al. (2020); Smith et al. (2021) only work for infinitesimal step size and the tools in Appendix E of Beneventano (2023) are the only ones which allow us to consider more products and big learning rate. Let $k=n/b$ be the number of batches in an epoch and call $B_{j}$ the batch of $j-th$ step and $\mathbf{a}_{B_{j}},\mathbf{b}_{B_{j}}$ the averages of $\mathbf{a}$ and $\mathbf{b}$ over the batch $B_{j}$.
@@
Here if $\eta=O\left(\frac{b^{2}}{n}[\mathrm{Var}(\mathbf{a}_{1})]^{-1}\right)$ which is generally very small, the second them may already be of size $O(1)$. We already see that in this case $\eta$ of the instability threshold has to be set also through the variance of the $\mathbf{a}_{i}$, the size of the batch or the number of batches. More generally, by applying the results in Appendix E of Beneventano (2023) we obtain that